OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 15 meters and legs that measure 9 meters and 12 meters. What is the surface area of this prism if the height of the prism is 11 meters?

OpenStudy (anonymous):

I'm not another student just looking for an answer. I'd really like to know how to do this. Please help!

OpenStudy (anonymous):



OpenStudy (anonymous):

find the base area = 1/2 * 9 * 12 *2 Find the sloping face area = 9*11 + 12*11 + 15*11 add it all u will get the answer

OpenStudy (anonymous):

If I did it correctly.. is it 504 meters squared?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Yeah, it's right! I just didn't understand how the figure would look set up, and where to go from there. Thanks, dawg. ;D